OpenStudy (anonymous):

a circumference of a circle of radius r is given by the polynomial 2(3.14)r. find circumference of a circle with a radius 9 cm

OpenStudy (mathstudent55):

C = 2(3.14)r Let r = 9 cm, and solve for C

OpenStudy (anonymous):

would it =56.52 or did i do it wrong im confused

OpenStudy (mathstudent55):

C = 2(3.14)(9 cm) = 6.28(9 cm) = 56.52 cm. You are correct.